what is the percentage yield of 9.28 g of oxygen react with hydrogen to produce 8.81 g of water? Show the solution

1
Expert's answer
2022-05-30T02:34:05-0400

2H2 + O2 = 2H2O

n(O2) = 9.28/32 = 0.29 mole

n(H2O) = 0.58 mole

m(H2O) = 10.44 g

yield = 8.81/10.44 =

84.4%